OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. Will give a medal

OpenStudy (anonymous):

Are you solving this or do you need to write it in a different form?

OpenStudy (anonymous):

I'm trying to solv it

OpenStudy (anonymous):

OK, so first.. Can you write the problem using symbols and numbers for me =)

OpenStudy (anonymous):

9 + n = 12 - 3n ??

OpenStudy (anonymous):

Is it 9 * n = 12 - 3 * n ?

OpenStudy (anonymous):

Product = Multiplication Anyways, you have to make it to where the x is only on one side of the equation. To start off, we add, sinc we are subtracting in the original equation, 3n to both sides of the equation 9n + 3n = 12 - 3n + 3n 9n + 3n = 12 Can you tell me what is 9n + 3n?

OpenStudy (anonymous):

12n

OpenStudy (anonymous):

Correct, so 12n = 12 So you know what to do next?

OpenStudy (anonymous):

I'm not sure what to do now

OpenStudy (anonymous):

You divide 12 on both sides of the equation to make n alone 12n/12 = 12/12 n = 12/12 and we know that 12/12 = 1 so n = 1 =)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Aaaaanndd \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) If you need anymore help tag me @MSxSammiBooxX or @sammixboo

OpenStudy (anonymous):

And if you need any help with using OpenStudy, private message me =3